OpenStudy (anonymous):

simplify 10/14 to lowest terms

OpenStudy (anonymous):

.714

OpenStudy (klimenkov):

\(\frac{10}{14}=\frac{2\cdot5}{2\cdot7}=\frac22\cdot\frac57\)